# Retrieval-Augmented Generation in Biomedicine: A Survey of Technologies, Datasets, and Clinical Applications

### v1

@@ Introduction
Type 4: LLM-based embedding. Generative LLMs have demonstrated remarkable generative capabilities across diverse NLP tasks, owing to their sophisticated understanding of language and context. With billions of parameters and training on vast corpora, these models develop rich internal representations that capture nuanced semantic relationships and contextual dependencies in text 247. Recognizing this powerful language comprehension ability, researchers have begun leveraging open-source LLMs as embedding generators for retrieval tasks, extracting representations from their internal layers to create semantically meaningful document and query embeddings that outperform traditional embedding approaches. As shown in Table LABEL:tab:Classification, notable implementations include BiomedRAG 128, which utilizes MedLLaMA 13b 224, SurgeryLLM 163, which employs GPT4All 16, and the comparative study by Myers et al. 160, which evaluates LLM2Vec 22 for clinical information retrieval applications. However, most LLMs are decoder-only architectures that employ causal attention mechanisms, which inherently limit their ability to generate rich contextualized representations. Thus, researchers are increasingly focusing on developing more advanced LLM-based embedding methods, such as Llama2Vec 138, Landmark Embedding 143, which provide promising dense retrieval alternatives in biomedical RAG.
@@
Type 5: Domain adaptation retrievers. Although the aforementioned off-the-shelf retrievers are readily available, the task of searching for relevant and accurate documents in the medical domain remains challenging. Consequently, customized retrievers tailored specifically to scientific and biomedical tasks have been developed. As shown in Table LABEL:tab:Classification, these advanced retrievers include Contriever 92, MedCPT 97, SPECTER 37, BMRetriever 234, and the MEDRAG toolkit 128. These specialized retrieval architectures demonstrate the importance of domain adaptation in biomedical information retrieval, significantly improving retrieval performance in several studies 93, 3, 242. For example, Jeong et al. 93 proposed Self-BioRAG, which leverages MedCPT to improve medical question answering through a self-augmentation approach. Similarly, the MEDRAG framework 6 combines multiple retrieval strategies, including BM25, SPECTER, Contriever, and MedCPT to create a comprehensive biomedical RAG tool. RAMIE utilizes MedCPT, Contriever, and BMRetriever as retrievers, builds a RAG framework to extract diverse types of information about dietary supplements (DSs) from clinical records, enhanced overall accuracy in its experiments 242. These implementations highlight how domain-adapted retrievers benefit biomedical RAG systems.
@@
Dense Retriever Infrastructure. Vector stores provide an infrastructure element for dense retrievers in semantic search systems 80. These specialized databases efficiently index, store, and search high-dimensional vector embeddings, offering optimized methods like K-nearest neighbors (KNN) for RAG developers. Vector databases enable rapid retrieval of semantically similar documents at scale, significantly enhancing the performance and response time of biomedical RAG systems. The process of embedding, storing, and searching documents involves transforming text into vector representations, organizing them in specialized data structures, and efficiently retrieving the most relevant matches. Based on our survey of biomedical RAG publications presented in Table LABEL:tab:Classification, we have compiled the most commonly used vector databases in Table 3. In Table LABEL:tab:Classification, it can be observed that the most popular vector stores are FAISS and Chroma, which were utilized 4 times and 3 times respectively.
@@
Chroma 36 AI-native open-source vector database with built-in functionality
Faiss 52 Library for efficient similarity search and clustering of dense vectors
Pinecone 172 Production-ready vector database for similarity search at scale
Weaviate 220 Graph-based vector search engine with semantic capabilities
HNSW 147 Hierarchical navigable small world graphs for approximate nearest neighbor search
@@
Knowledge graph retrievers offer significant advantages in biomedical RAG by providing structured, interlinked domain knowledge that overcomes several limitations of traditional retrieval methods. Unlike vector-based approaches, knowledge graphs capture explicit relationships between biomedical entities, enabling more precise retrieval of contextually relevant information, enhanced reasoning capabilities, and improved explainability171. Table 4 presents knowledge graph-based biomedical RAG studies published between 2024 and 2025, outlining each system’s application domains, and knowledge graph source. As we can notice, the application areas vary significantly, including medical question answering 194, knowledge extraction 47, and diagnosis prediction 64.
@@
KG-RAG 194 Biomedical text generation, medical question answering SPOKE biomedical knowledge graph 2024
HEALIE 108 Personalized medical content generation HEALIE KG (customized) 2024
KRAGEN 152 Biomedical problem solving Alzheimer’s knowledge graph (AlzKB) 2024
Ascle 237 Various medical text generation tasks UMLS 2024
DALK 125 Alzheimer’s QA Alzheimer’s knowledge graph (AlzKB) 2024
Gilbert et al. 69 Medical information structuring and interlinking Not specified 2024
NetMe 2.0 47 Biomedical knowledge extraction BKGs constructed using OntoTagMe and Wikidata 2024
NEKO 229 Knowledge mining in synthetic biology PubMed-derived knowledge graphs 2025
DR.KNOWS 64 Diagnosis prediction from EHRs UMLS 2025
ESCARGOT 151 Biomedical reasoning and knowledge retrieval Alzheimer’s knowledge graph (AlzKB) 2025
@@
  • Biomedical knowledge graph construction or curation involves the systematic process of building, updating, or adapting knowledge graphs from domain-specific ontologies, literature, and other data sources to create structured biomedical knowledge repositories. NetMe 2.0 47 constructs biomedical knowledge graphs on-the-fly with OntoTagMe for entity linking while NEKO 229 mines PubMed literature for synthetic biology knowledge graphs. Differently, KG-RAG 194 utilizes the pre-existing SPOKE biomedical knowledge graph.
  • Query to knowledge graph transformation encompasses the conversion of natural language queries or user inputs into structured queries that efficiently retrieve relevant subgraphs from the knowledge graph. ESCARGOT 151 employs Cypher queries for precise retrieval, KRAGEN 152 vectorizes knowledge graphs for semantic searching, and DR.KNOWS 64 transforms queries to retrieve UMLS-based diagnostic pathways aligned with patient data.
  • Path-based or entity-centric knowledge extraction refers to the process of extracting relevant knowledge patterns, paths, or entity-centered information from the knowledge graph based on the transformed query. DR.KNOWS 64 utilizes attention-based path ranking for diagnostic reasoning, KRAGEN 152 applies graph-of-thoughts for multi-hop relationships, and HEALIE 108 extracts paths incorporating both medical knowledge and socio-economic factors.
  • Strategic integration into LLM prompts involves methods for incorporating retrieved knowledge into prompts for LLMs in ways that maximize factual accuracy while optimizing token usage and computational efficiency. KG-RAG 194 implements token-optimized retrieval to reduce consumption, ESCARGOT 151 selectively integrates knowledge through dynamic Graph of Thoughts, and NetMe 2.0 47 transforms graph data into natural language prompts via its Graph-RAG module.
@@
Multimodal biomedical RAG represents an advanced approach to healthcare AI that integrates LLMs with multiple types of data inputs beyond text, such as medical images, electrocardiogram (ECG), and other clinical measurements. These systems can enhance diagnostic accuracy and clinical decision support by leveraging the complementary strengths of different data modalities, retrieving relevant information from specialized knowledge bases, and generating comprehensive insights that would be impossible to achieve with single-modality approaches 4. As shown in Table 5, multimodal biomedical RAG systems have been successfully deployed for amblyopia detection 215, prescription label interpretation 208, lung cancer staging 211, digital pathology analysis 162, and automated radiology report generation 178, 199, 225, 226, 238.
@@
Ranjit et al. 178 Chest X-Ray report generation Chest X-ray and text 2023
FactMM-RAG 199 Radiology report generation X-Ray and text 2024
MMed-RAG 225 Medical VQA & Report generation Medical images and text 2024
MEAG 215 Amblyopia diagnosis Eye tracking and text 2024
Tozuka et al.  211 Lung cancer staging CT findings and TNM classifications 2024
GPT4DFCI-RAG 162 Digital pathology QA Pathology images and text 2024
Raminedi et al. 177 Radiology report generation X-ray images and text 2024
RULE 226 Medical VQA & Report generation Medical images and text 2024
Thetbanthad et al. 208 Prescription label identification Image (labels) and text 2025
Hu et al. 87 Pathology report generation Whole slide images and text 2025
STREAM 238 Chest X-ray report generation Chest X-ray and text 2025
Figure 3 illustrates the architectural workflow of a multimodal biomedical RAG system. The key distinction between multimodal biomedical RAG and traditional biomedical RAG lies in data input diversity and processing architecture. While traditional RAG systems primarily process textual queries to retrieve text-based medical knowledge, multimodal RAG incorporates and processes non-textual data (images, sensor readings, etc.) alongside text, requiring specialized encoders for each modality to meaningfully integrate these diverse inputs 4. The basic framework of multimodal biomedical RAG typically consists of: (1) modality-specific encoders (e.g., Vision Transformers 32 for images, specialized encoders for time-series data 32); (2) knowledge retrieval components that access domain-specific databases based on multimodal query representations; and (3) generation modules that synthesize comprehensive outputs informed by both the multimodal inputs and retrieved knowledge. For instance, Ranjit et al.178, FactMM-RAG 199 and STREAM 238 use multimodal embeddings for retrieval of relevant candidate radiology text to generate Chest X-Ray report. MMed-RAG 225 and RULE 226 build a multimodal RAG system for Medical Visual Question Answering (Medical VQA) and report generation. The MEAG 215 system uses specialized encoders for eye movement data alongside text, Raminedi et al. 162 employ Vision Transformers as encoders for radiological images with cross-attention mechanisms for fusion with text, while Thetbanthad et al. 208 first converts prescription label images to text through OCR before retrieving relevant pharmaceutical information from DrugBank.
@@
Rerankers in biomedical RAG systems serve as intermediary components that refine and prioritize the relevance of initially retrieved documents before they are processed by the generative models 190. Positioned between retrievers and generators, rerankers apply sophisticated algorithms to re-score and re-order candidate documents based on their contextual relevance to the query, effectively functioning as a second-stage filtering mechanism that enhances the precision of information ultimately provided to the generation component. By prioritizing the most pertinent information, rerankers help to minimize hallucinations and improve the veracity of generated outputs.
Biomedical RAG rerankers can be categorized into distinct types based on their underlying methodologies. Statistical rerankers apply traditional information retrieval techniques; for instance, CliniqIR 3 and MEDRAG 6 both implement Reciprocal Rank Fusion (RRF) to combine multiple retrieval signals, while DRAGON-AI 209 employs Maximal Marginal Relevance (MMR) to balance relevance with diversity. Content-based rerankers, in contrast, focus on content similarity; notably, WeiseEule 6 introduces a keyword frequency-based ranking system specifically designed for biomedical literature retrieval. Model-based rerankers employ transformer architectures to perform contextual relevance assessment; specifically, PodGPT 94 uses a general embedding model (bge-reranker-large) for scientific literature retrieval, whereas the CHA2DS2-VASc risk factor extraction system 5 utilizes a specialized cross-encoder model known as BGE M3-Embedding 31 to enhance clinical document relevance scoring. Within this category, some domain adaptation retrievers incorporate a reranker as a component of their model; for example, MedCPT 97 trains a specific reranker denoted as CrossEnc. Additionally, LLM-based rerankers leverage foundation models’ reasoning capabilities; GNQA 41 employs GPT-3.5/4/4o as zero-shot binary classifiers for relevance determination, and LmRaC 39 implements an LLM-based paragraph usefulness assessment with a minimum threshold score of 7. (on a 10 point scale). Finally, hybrid approaches combine multiple strategies to enhance performance. For example, Clinfo.ai 141 combines LLM classification with BM25 to re-rank retrieved articles.
@@
The generation component represents the last element in biomedical RAG systems, responsible for synthesizing retrieved information into human-like responses that address user queries. In biomedical contexts, this process requires balancing natural language fluency with domain-specific accuracy, as generated content often informs healthcare decisions 122. This section examines various generation models employed in recent biomedical RAG studies.
@@
As shown in Table LABEL:tab:llm_taxonomy, the majority of LLMs employ decoder-only architectures, with T5 being a notable exception. This architectural preference stems from several inherent advantages for generative tasks, particularly the natural suitability for next-token prediction and significantly improved parameter efficiency 247. Among open-source LLMs, LLaMa 2 210 is the most popular, with 15 biomedical applications built upon it. LLaMa 3 7, Mistral 7B 155, and T5 176 are also widely used. These open-source LLMs have various sizes, from 0.06B (T5) to 671B (Deepseek R1), providing flexible options for building Biomedical RAG systems. Recently, several powerful general open-source LLMs have emerged, including, Mistral Small 3.1 156 and Gemma 3 72. Although we have not found these newer models being used for biomedical RAG systems in our survey, we believe they show promise due to their outstanding performance in general domain tasks.
For commercial LLMs, the ChatGPT series APIs maintain a dominant position, presumably attributable to their superior performance. Nevertheless, several studies have demonstrated that Claude and Gemini achieve comparable efficacy when implemented in biomedical RAG applications 217, 84, 119, thereby offering viable alternatives to ChatGPT. Furthermore, it is noteworthy that both DeepSeek and Mistral have established dual accessibility paradigms: providing open-source parameters for their base models while simultaneously offering commercial API services.
In contrast to open-source and commercial LLMs, biomedical specialized LLMs have been relatively underutilized in biomedical RAG applications. Recent studies 166, 242, 11 have demonstrated that within RAG frameworks, biomedical specialized LLMs do not consistently outperform open-source alternatives with statistical significance. Thus, the more effective methodologies for leveraging biomedical specialized LLMs within retrieval-augmented generation frameworks remain an area requiring substantial further investigation.
@@
T5 176 0.06-11B Encoder-Decoder VAIV 115, CLEAR 140, DR.KNOWS 64
ChatGLM3 70 6B Decoder-only COPD 244, ChatZOC 144
LLaMA 2 210 7-70B Decoder-only Guo et al. 76, KREIMEYER et al. 117, CaLM 168, MMRAG 241, Yu et al. 239, RAMIE 242, EyeGPT 33, KG-RAG 194, JMLR 219, BiomedRAG 128, Self-BioRAG 93, CaLM 168, Du et al. 49, SurgeryLLM 163, Kreimeyer et al. 117
LLaMA 3 7 8-70B Decoder-only MMRAG 241, Fatharanihttps et al. 54, RAMIE 242, Woo et al. 222, RAG-HPO 67, i-MedRAG 232, Hewitt et al. 84, PodGPT 94, SurgeryLLM 163
Phi-3 Mini 154 3.8B Decoder-only Fatharanihttps et al. 54
Mistral 7B 155 7B Decoder-only Boulos et al. 110, RAMIE 242, Woo et al. 222, RAGPR 248, LITURAt 169, Kreimeyer et al. 117
QWen 175 32B Decoder-only NEKO 229, Thetbanthad 208
Baichuan 216 7-13B Decoder-only ChatZOC 144
Falcon 12 7-180B Decoder-only AskFDALabel 19, CaLM 168
Zephyr 213 7B Decoder-only Moser et al. 159
AceGPT 89 7-13B Decoder-only InfectA-Chat 186
Gemma 62 1-27B Decoder-only PodGPT 94
Deepseek R1 44 671B Decoder-only Feng et al. 55
@@
ChatGPT-3.5/4/4o 165 Proprietary Decoder-only Woo et al. 222, RAGPR 248, ChatZOC 144, KG-RAG 194, Clinfo.ai 141, Yu et al. 239, ChatENT 139, CaLM 168, Rau et al. 180, Endo-chat 59, BiomedRAG 128, Lu 142, Tan 15, GNQA 40, Puts 173, DRAGON-AI 209, Choi 35, RISE 217, Du et al. 49,i-MedRAG 232 , Hewitt et al. 84, RECTIFIER 214, Kainer et al. 107, Gong et al. 71, VAIA 115, RareDxGPT 240, Selcuk et al. 186, FAVOR-GPT 130, Markey et al. 148,
Claude-3.5 18 Proprietary Decoder-only Woo et al. 222, RISE 217, Hewitt et al. 84
Gemini 73 Proprietary Multimodal Upadhyaya et al. 215, Tozuka et al. 211, MEREDITH 119
@@
MedAlpaca 79 7-13B LLaMA-based RAMIE 242
PMC-LLaMA 223 7B LLaMA-based RAMIE 242, BiomedRAG 128, Ozaki et al. 166
BioMistral 118 7B Mistral-based RAMIE 242, pRAGe 28
MEDITRON 34 7-70B LLaMA-based Ozaki et al. 166, Alkaeed et al. 11
Meerkat 113 7B Mistral-based Sohn et al. 192
@@
Domain adaptation in biomedical RAG systems constitutes a methodological framework that can be strategically implemented across various RAG pipeline stages to optimize performance for medical applications 189. This approach facilitates enhanced performance on biomedical tasks while simultaneously preserving the fundamental capabilities inherent to foundation models. Based on our systematic analysis of the literature, we propose a taxonomy of domain adaptation techniques categorized according to their implementation within specific RAG pipeline components.
@@
Retriever Stage Adaptation. Specialized retrieval components constitute the initial layer of domain adaptation, wherein techniques such as contrastive learning (implemented in Contriever 92 and MedCPT 97) enable the development of more semantically rich biomedical embeddings. These adaptations augment the system’s capacity to identify and retrieve pertinent medical documents through enhanced comprehension of domain-specific terminology and conceptual relationships. Additional retriever adaptation methodologies include continued pre-training on biomedical corpora (as demonstrated in BMRETRIEVER 234) and domain-specific architectural modifications (exemplified by LADER 98 with PubMedBERT initialization). These approaches substantially enhance retrieval precision for specialized medical literature, effectively mitigating the lexical and semantic disparities between general and biomedical language domains.

### v2

@@ Introduction
Type 4: LLM-based embedding. Generative LLMs have demonstrated remarkable generative capabilities across diverse NLP tasks, owing to their sophisticated understanding of language and context. With billions of parameters and training on vast corpora, these models develop rich internal representations that capture nuanced semantic relationships and contextual dependencies in text 12. Recognizing this powerful language comprehension ability, researchers have begun leveraging open-source LLMs as embedding generators for retrieval tasks, extracting representations from their internal layers to create semantically meaningful document and query embeddings that outperform traditional embedding approaches. As shown in Table LABEL:tab:Classification, notable implementations include BiomedRAG 4, which utilizes MedLLaMA 13b 111, SurgeryLLM 28, which employs GPT4All 103, and the comparative study by Myers et al. 104, which evaluates LLM2Vec 105 for clinical information retrieval applications. However, most LLMs are decoder-only architectures that employ causal attention mechanisms, which inherently limit their ability to generate rich contextualized representations. Thus, researchers are increasingly focusing on developing more advanced LLM-based embedding methods, such as Llama2Vec 112, Landmark Embedding 113, which provide promising dense retrieval alternatives in biomedical RAG.
@@
Type 5: Domain adaptation retrievers. Although the aforementioned off-the-shelf retrievers are readily available, the task of searching for relevant and accurate documents in the medical domain remains challenging. Consequently, customized retrievers tailored specifically to scientific and biomedical tasks have been developed. As shown in Table LABEL:tab:Classification, these advanced retrievers include Contriever 57, MedCPT 33, SPECTER 22, BMRetriever 55, and the MEDRAG toolkit 4. These specialized retrieval architectures demonstrate the importance of domain adaptation in biomedical information retrieval, significantly improving retrieval performance in several studies 106, 52, 107. For example, Jeong et al. 106 proposed Self-BioRAG, which leverages MedCPT to improve medical question answering through a self-augmentation approach. Similarly, the MEDRAG framework 101 combines multiple retrieval strategies, including BM25, SPECTER, Contriever, and MedCPT to create a comprehensive biomedical RAG tool. RAMIE utilizes MedCPT, Contriever, and BMRetriever as retrievers, builds a RAG framework to extract diverse types of information about dietary supplements (DSs) from clinical records, enhanced overall accuracy in its experiments 107. These implementations highlight how domain-adapted retrievers benefit biomedical RAG systems.
@@
Dense Retriever Infrastructure. Vector stores provide an infrastructure element for dense retrievers in semantic search systems 114. These specialized databases efficiently index, store, and search high-dimensional vector embeddings, offering optimized methods like K-nearest neighbors (KNN) for RAG developers. Vector databases enable rapid retrieval of semantically similar documents at scale, significantly enhancing the performance and response time of biomedical RAG systems. The process of embedding, storing, and searching documents involves transforming text into vector representations, organizing them in specialized data structures, and efficiently retrieving the most relevant matches. Based on our survey of biomedical RAG publications presented in Table LABEL:tab:Classification, we have compiled the most commonly used vector databases in Table 3. In Table LABEL:tab:Classification, it can be observed that the most popular vector stores are FAISS and Chroma, which were utilized 4 times and 3 times respectively.
@@
Chroma 115 AI-native open-source vector database with built-in functionality
Faiss 116 Library for efficient similarity search and clustering of dense vectors
Pinecone 117 Production-ready vector database for similarity search at scale
Weaviate 118 Graph-based vector search engine with semantic capabilities
HNSW 78 Hierarchical navigable small world graphs for approximate nearest neighbor search
@@
Knowledge graph retrievers offer significant advantages in biomedical RAG by providing structured, interlinked domain knowledge that overcomes several limitations of traditional retrieval methods. Unlike vector-based approaches, knowledge graphs capture explicit relationships between biomedical entities, enabling more precise retrieval of contextually relevant information, enhanced reasoning capabilities, and improved explainability119. Table 4 presents knowledge graph-based biomedical RAG studies published between 2024 and 2025, outlining each system’s application domains, and knowledge graph source. As we can notice, the application areas vary significantly, including medical question answering 100, knowledge extraction 120, and diagnosis prediction 17.
@@
KG-RAG 100 Biomedical text generation, medical question answering SPOKE biomedical knowledge graph 2024
HEALIE 29 Personalized medical content generation HEALIE KG (customized) 2024
KRAGEN 121 Biomedical problem solving Alzheimer’s knowledge graph (AlzKB) 2024
Ascle 30 Various medical text generation tasks UMLS 2024
DALK 122 Alzheimer’s QA Alzheimer’s knowledge graph (AlzKB) 2024
Gilbert et al. 123 Medical information structuring and interlinking Not specified 2024
NetMe 2.0 120 Biomedical knowledge extraction BKGs constructed using OntoTagMe and Wikidata 2024
NEKO 124 Knowledge mining in synthetic biology PubMed-derived knowledge graphs 2025
DR.KNOWS 17 Diagnosis prediction from EHRs UMLS 2025
ESCARGOT 125 Biomedical reasoning and knowledge retrieval Alzheimer’s knowledge graph (AlzKB) 2025
@@
  • Biomedical knowledge graph construction or curation involves the systematic process of building, updating, or adapting knowledge graphs from domain-specific ontologies, literature, and other data sources to create structured biomedical knowledge repositories. NetMe 2.0 120 constructs biomedical knowledge graphs on-the-fly with OntoTagMe for entity linking while NEKO 124 mines PubMed literature for synthetic biology knowledge graphs. Differently, KG-RAG 100 utilizes the pre-existing SPOKE biomedical knowledge graph.
  • Query structuration encompasses the conversion of natural language queries or user inputs into structured queries that efficiently retrieve relevant subgraphs from the knowledge graph. ESCARGOT 125 employs Cypher queries for precise retrieval, KRAGEN 121 vectorizes knowledge graphs for semantic searching, and DR.KNOWS 17 transforms queries to retrieve UMLS-based diagnostic pathways aligned with patient data.
  • Path-based or entity-centric knowledge extraction refers to the process of extracting relevant knowledge patterns, paths, or entity-centered information from the knowledge graph based on the transformed query. DR.KNOWS 17 utilizes attention-based path ranking for diagnostic reasoning, KRAGEN 121 applies graph-of-thoughts for multi-hop relationships, and HEALIE 29 extracts paths incorporating both medical knowledge and socio-economic factors.
  • Strategic integration into LLM prompts involves methods for incorporating retrieved knowledge into prompts for LLMs in ways that maximize factual accuracy while optimizing token usage and computational efficiency. KG-RAG 100 implements token-optimized retrieval to reduce consumption, ESCARGOT 125 selectively integrates knowledge through dynamic Graph of Thoughts, and NetMe 2.0 120 transforms graph data into natural language prompts via its Graph-RAG module.
@@
Multimodal biomedical RAG represents an advanced approach to healthcare AI that integrates LLMs with multiple types of data inputs beyond text, such as medical images, electrocardiogram (ECG), and other clinical measurements. These systems can enhance diagnostic accuracy and clinical decision support by leveraging the complementary strengths of different data modalities, retrieving relevant information from specialized knowledge bases, and generating comprehensive insights that would be impossible to achieve with single-modality approaches 126. As shown in Table 5, multimodal biomedical RAG systems have been successfully deployed for amblyopia detection 32, prescription label interpretation 127, lung cancer staging 128, digital pathology analysis 129, and automated radiology report generation 130, 31, 131, 132, 133.
@@
Ranjit et al. 130 Chest X-Ray report generation Chest X-ray and text 2023
FactMM-RAG 31 Radiology report generation X-Ray and text 2024
MMed-RAG 131 Medical VQA & Report generation Medical images and text 2024
MEAG 32 Amblyopia diagnosis Eye tracking and text 2024
Tozuka et al.  128 Lung cancer staging CT findings and TNM classifications 2024
GPT4DFCI-RAG 129 Digital pathology QA Pathology images and text 2024
Raminedi et al. 19 Radiology report generation X-ray images and text 2024
RULE 132 Medical VQA & Report generation Medical images and text 2024
Thetbanthad et al. 127 Prescription label identification Image (labels) and text 2025
Hu et al. 134 Pathology report generation Whole slide images and text 2025
STREAM 133 Chest X-ray report generation Chest X-ray and text 2025
Figure 3 illustrates the architectural workflow of a multimodal biomedical RAG system. The key distinction between multimodal biomedical RAG and traditional biomedical RAG lies in data input diversity and processing architecture. While traditional RAG systems primarily process textual queries to retrieve text-based medical knowledge, multimodal RAG incorporates and processes non-textual data (images, sensor readings, etc.) alongside text, requiring specialized encoders for each modality to meaningfully integrate these diverse inputs 126. The basic framework of multimodal biomedical RAG typically consists of: (1) modality-specific encoders (e.g., Vision Transformers 135 for images, specialized encoders for time-series data 135); (2) knowledge retrieval components that access domain-specific databases based on multimodal query representations; and (3) generation modules that synthesize comprehensive outputs informed by both the multimodal inputs and retrieved knowledge. For instance, Ranjit et al.130, FactMM-RAG 31 and STREAM 133 use multimodal embeddings for retrieval of relevant candidate radiology text to generate Chest X-Ray report. MMed-RAG 131 and RULE 132 build a multimodal RAG system for Medical Visual Question Answering (Medical VQA) and report generation. The MEAG 32 system uses specialized encoders for eye movement data alongside text, Raminedi et al. 129 employ Vision Transformers as encoders for radiological images with cross-attention mechanisms for fusion with text, while Thetbanthad et al. 127 first converts prescription label images to text through OCR before retrieving relevant pharmaceutical information from DrugBank.
@@
Rerankers in biomedical RAG systems serve as intermediary components that refine and prioritize the relevance of initially retrieved documents before they are processed by the generative models 45. Positioned between retrievers and generators, rerankers apply sophisticated algorithms to re-score and re-order candidate documents based on their contextual relevance to the query, effectively functioning as a second-stage filtering mechanism that enhances the precision of information ultimately provided to the generation component. By prioritizing the most pertinent information, rerankers help to minimize hallucinations and improve the veracity of generated outputs.
Biomedical RAG rerankers can be categorized into distinct types based on their underlying methodologies. Statistical rerankers apply traditional information retrieval techniques; for instance, CliniqIR 52 and MEDRAG 101 both implement Reciprocal Rank Fusion (RRF) to combine multiple retrieval signals, while DRAGON-AI 83 employs Maximal Marginal Relevance (MMR) to balance relevance with diversity. Content-based rerankers, in contrast, focus on content similarity; notably, WeiseEule 101 introduces a keyword frequency-based ranking system specifically designed for biomedical literature retrieval. Model-based rerankers employ transformer architectures to perform contextual relevance assessment; specifically, PodGPT 34 uses a general embedding model (bge-reranker-large) for scientific literature retrieval, whereas the CHA2DS2-VASc risk factor extraction system 136 utilizes a specialized cross-encoder model known as BGE M3-Embedding 137 to enhance clinical document relevance scoring. Within this category, some domain adaptation retrievers incorporate a reranker as a component of their model; for example, MedCPT 33 trains a specific reranker denoted as CrossEnc. Additionally, LLM-based rerankers leverage foundation models’ reasoning capabilities; GNQA 77 employs GPT-3.5/4/4o as zero-shot binary classifiers for relevance determination, and LmRaC 138 implements an LLM-based paragraph usefulness assessment with a minimum threshold score of 7. (on a 10 point scale). Finally, hybrid approaches combine multiple strategies to enhance performance. For example, Clinfo.ai 14 combines LLM classification with BM25 to re-rank retrieved articles.
@@
The generation component represents the last element in biomedical RAG systems, responsible for synthesizing retrieved information into human-like responses that address user queries. In biomedical contexts, this process requires balancing natural language fluency with domain-specific accuracy, as generated content often informs healthcare decisions 139. This section examines various generation models employed in recent biomedical RAG studies.
@@
As shown in Table LABEL:tab:llm_taxonomy, the majority of LLMs employ decoder-only architectures, with T5 being a notable exception. This architectural preference stems from several inherent advantages for generative tasks, particularly the natural suitability for next-token prediction and significantly improved parameter efficiency 12. Among open-source LLMs, LLaMa 2 140 is the most popular, with 15 biomedical applications built upon it. LLaMa 3 38, Mistral 7B 141, and T5 142 are also widely used. These open-source LLMs have various sizes, from 0.06B (T5) to 671B (Deepseek R1), providing flexible options for building Biomedical RAG systems. Recently, several powerful general open-source LLMs have emerged, including, Mistral Small 3.1 40 and Gemma 3 143. Although we have not found these newer models being used for biomedical RAG systems in our survey, we believe they show promise due to their outstanding performance in general domain tasks.
For commercial LLMs, the ChatGPT series APIs maintain a dominant position, presumably attributable to their superior performance. Nevertheless, several studies have demonstrated that Claude and Gemini achieve comparable efficacy when implemented in biomedical RAG applications 85, 144, 15, thereby offering viable alternatives to ChatGPT. Furthermore, it is noteworthy that both DeepSeek and Mistral have established dual accessibility paradigms: providing open-source parameters for their base models while simultaneously offering commercial API services.
In contrast to open-source and commercial LLMs, biomedical specialized LLMs have been relatively underutilized in biomedical RAG applications. Recent studies 145, 107, 146 have demonstrated that within RAG frameworks, biomedical specialized LLMs do not consistently outperform open-source alternatives with statistical significance. Thus, the more effective methodologies for leveraging biomedical specialized LLMs within retrieval-augmented generation frameworks remain an area requiring substantial further investigation.
@@
T5 142 0.06-11B Encoder-Decoder VAIV 53, CLEAR 99, DR.KNOWS 17
ChatGLM3 147 6B Decoder-only COPD 148, ChatZOC 149
LLaMA 2 140 7-70B Decoder-only Guo et al. 27, KREIMEYER et al. 150, CaLM 26, MMRAG 151, Yu et al. 152, RAMIE 107, EyeGPT 153, KG-RAG 100, JMLR 3, BiomedRAG 4, Self-BioRAG 106, CaLM 26, Du et al. 154, SurgeryLLM 28, Kreimeyer et al. 150
LLaMA 3 38 8-70B Decoder-only MMRAG 151, Fatharanihttps et al. 155, RAMIE 107, Woo et al. 156, RAG-HPO 72, i-MedRAG 157, Hewitt et al. 144, PodGPT 34, SurgeryLLM 28
Phi-3 Mini 158 3.8B Decoder-only Fatharanihttps et al. 155
Mistral 7B 141 7B Decoder-only Boulos et al. 159, RAMIE 107, Woo et al. 156, RAGPR 80, LITURAt 62, Kreimeyer et al. 150
QWen 160 32B Decoder-only NEKO 124, Thetbanthad 127
Baichuan 161 7-13B Decoder-only ChatZOC 149
Falcon 162 7-180B Decoder-only AskFDALabel 42, CaLM 26
Zephyr 163 7B Decoder-only Moser et al. 164
AceGPT 165 7-13B Decoder-only InfectA-Chat 94
Gemma 166 1-27B Decoder-only PodGPT 34
Deepseek R1 41 671B Decoder-only Feng et al. 167
@@
ChatGPT-3.5/4/4o 35 Proprietary Decoder-only Woo et al. 156, RAGPR 80, ChatZOC 149, KG-RAG 100, Clinfo.ai 14, Yu et al. 152, ChatENT 89, CaLM 26, Rau et al. 168, Endo-chat 88, BiomedRAG 4, Lu 18, Tan 169, GNQA 77, Puts 170, DRAGON-AI 83, Choi 171, RISE 85, Du et al. 154,i-MedRAG 157 , Hewitt et al. 144, RECTIFIER 91, Kainer et al. 95, Gong et al. 49, VAIA 53, RareDxGPT 16, Selcuk et al. 94, FAVOR-GPT 87, Markey et al. 172,
Claude-3.5 36 Proprietary Decoder-only Woo et al. 156, RISE 85, Hewitt et al. 144
Gemini 37 Proprietary Multimodal Upadhyaya et al. 32, Tozuka et al. 128, MEREDITH 15
@@
MedAlpaca 173 7-13B LLaMA-based RAMIE 107
PMC-LLaMA 174 7B LLaMA-based RAMIE 107, BiomedRAG 4, Ozaki et al. 145
BioMistral 175 7B Mistral-based RAMIE 107, pRAGe 176
MEDITRON 177 7-70B LLaMA-based Ozaki et al. 145, Alkaeed et al. 146
Meerkat 178 7B Mistral-based Sohn et al. 179
@@
Domain adaptation in biomedical RAG systems constitutes a methodological framework that can be strategically implemented across various RAG pipeline stages to optimize performance for medical applications 180. This approach facilitates enhanced performance on biomedical tasks while simultaneously preserving the fundamental capabilities inherent to foundation models. Based on our systematic analysis of the literature, we propose a taxonomy of domain adaptation techniques categorized according to their implementation within specific RAG pipeline components.
@@
Retriever Stage Adaptation. Specialized retrieval components constitute the initial layer of domain adaptation, wherein techniques such as contrastive learning (implemented in Contriever 57 and MedCPT 33) enable the development of more semantically rich biomedical embeddings. These adaptations augment the system’s capacity to identify and retrieve pertinent medical documents through enhanced comprehension of domain-specific terminology and conceptual relationships. Additional retriever adaptation methodologies include continued pre-training on biomedical corpora (as demonstrated in BMRETRIEVER 55) and domain-specific architectural modifications (exemplified by LADER 97 with PubMedBERT initialization). These approaches substantially enhance retrieval precision for specialized medical literature, effectively mitigating the lexical and semantic disparities between general and biomedical language domains.